OpenStudy (anonymous):

Xplain how the graph of y2 is related to the graph of y1

OpenStudy (anonymous):

Y1=1/x-1(+2) y2=1/x+1(-2)

OpenStudy (e.mccormick):

Have you graphed them or at lest found some points like where they intersect and the places they cross the axis?

OpenStudy (e.mccormick):

If you aim for those things, you will find some relationships between the graphs. This should lead you to an answer. The fact that 1/x is involved in both of them is another huge clue, because it means x can never be 0.